# MnLargeSymbols’164 MnLargeSymbols’171

### v2

@@ Abstract
We study a family of mean field games arising in modeling the behavior of strategic economic agents which move across space maximizing their utility from consumption and have the possibility to accumulate resources for production (such as human capital). The resulting mean field game PDE system is not covered in the actual literature on the topic as it displays weaker assumptions on the regularity of the data (in particular Lipschitz continuity and boundedness of the objective are lost), state constraints, and a non-standard interaction term. We obtain a first result on the existence of solution of the mean field game PDE system.
@@
Giovanni Zanco has been supported by the F-NF grant “MFGPRE: Mean-field games: regularity, path-dependency, economics”.
@@ 1 Introduction
The departure point of the present paper is the problem of understanding the time-space evolution of an economic system, which is a crucial topic for economists and policy makers. Many papers have studied this problem under the assumption that decision about the system are taken by a unique agent (the so-called ”social planner”, which can be seen as the government of a nation): in this case the resulting mathematical problem is the optimal control of a Partial Differential Equation (PDE from now on), possibly stochastic (see e.g. [5, 6, 7, 19]).
@@
From the mathematical viewpoint this kind of problems can be studied using the methodological tools of mean field games (MFGs from now on), a theory introduced in the last 20 years (starting from the seminal papers of Lasry-Lions [22] and Huang-Caines-Malhamé [20]) which allow to shed some light on the outcome of such complex dynamics.
This theory have been successfully employed to study the behavior of some economic models where many small homogeneous agents interact, see e.g. [1]. However no paper tried to consider the case of the time-space evolution of economic variables, maybe as it displays many technical difficulties to which we return below in this introduction. [^4] We mention here the paper [18] where a kind of mean field approach is used but without assuming that the small agents are forward-looking.
Hence, the purpose of this paper is to study, from the mathematical point of view, a family of MFGs arising in modeling how the strategic interactions of small agents can shape the time-space evolution of economic variables.
@@
We assume, as a starting point, that the small agents are homogeneous, move across space maximizing their intertemporal utility from consumption minus a disutility from displacement, and have the possibility to accumulate resources for production (such as human capital). Hence their state variables, at any time $t\geq 0$, are the position $x(t)$ and the human capital $h(t)$. The control variables are the velocity $v(t)$ with which they move in the space, and the allocation of human capital production $s(t)$.
@@
The resulting MFG, even in the simplest case, turns out to be very difficult and quite far from what is covered in the actual literature on the topic (see e.g. the books [10, 9]). The main reasons are the presence of non standard interaction terms, the weaker regularity assumptions on the data (in particular Lipschitz continuity and boundedness of the objective are lost), the state constraints. In particular, concerning the interaction term (which we will denote by $F$), we observe that it takes into account the fact that the spatial aggregation of agents increases the (human) capital $h$, through the so-called spillover effects. In particular, because of such structure, this term in not globally Lipschitz with respect to the distribution. Moreover, the positivity state constraints on the human capital brings the need of treating carefully the HJB and the FP equation proving suitable estimates near the border where $h=0$.
We study such MFG using the PDE approach, i.e. we consider the associated PDE system of Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck [^5] Some work on Fokker-Planck equations for uncontrolled interacting systems with spatial structure and non-standard interaction term has been carried out in [16] and [29].(FP) equations. Our main result is Theorem 4.4 on the existence of solutions of such PDE system. This is a departure point to understand the equilibria of the economic system, which will be the object of future research.
@@ 2.1 The economic model
The state equations of each agent on time intervals $[t_{0},T]$ are as follows. The dynamics for the position is given by
@@
$f\colon\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}\quad\hbox{ is Lipschitz and increasing and such that $f(0)=0$.}$
@@
[…]
over all admissible control process $(v,s)\in\mathcal{K}$. Here $\rho>0$ is the intertemporal discount factor while $a\colon K\rightarrow\mathbb{R}$ is a given strictly convex cost function satisfying $a(0)=0$. We make the following further assumptions:
@@ 2.2 The Mean Field Game system
On $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ we consider the forward-backward system of PDEs
$\begin{cases}-\partial_{t}V(t,x,h)+\rho V(t,x,h)=H_{0}(p)+H_{1}(x,h,\mu(t),DV(t,x,h))+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V(t,x,h)+\frac{1}{2}\chi^{2}h^{2}D^{2}_{hh}V(t,x,h),\\ \partial_{t}\mu(t)=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0},\quad V(T,x,h)=0\quad\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (6)
whose solution is a couple $(V,\mu)$, where $V$ is a real-valued function on $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ and $\mu$ is a function on $[0,T]$ taking values in $\mathcal{P}_{2}$ (a precise notion of solution will be given in Definition 2.3 below).
@@
    $V$ is continuous on $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$;
@@
    $D_{x}V$ and $D_{h}V$ are defined on $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$;
@@
    ×$\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)$ $\mu(t;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(0,x,h)\mu_{0}(\mathrm{d}x,\mathrm{d}h)$ (7)
@@
    $+\left.D_{q}H_{1}\left(x,h,\mu(r),D_{h}V(r,x,h)\right)D_{h}\phi(r,x,h)\right)\mu(r;\mathrm{d}x,\mathrm{d}h)\mathrm{d}r$
    $+\frac{1}{2}\int_{0}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\left(\epsilon^{2}D^{2}_{xx}\phi(r,x,h)+\chi^{2}h^{2}D^{2}_{hh}\phi(r,x,h)\right)\mu(r;\mathrm{d}x,\mathrm{d}h)\mathrm{d}r$
    $+\int_{0}^{t}\int_{\mathbb{R}\times\mathbb{R}_{+}}\partial_{t}\phi(r,x,h)\mu(r;\mathrm{d}x,\mathrm{d}h)\mathrm{d}r.$
@@
In particular we will often deal with maps $[0,T]\ni t\mapsto\mu(t)$; in this case we write
@@ 3.1 The Hamilton-Jacobi-Bellman equation
To study the HJB equation we fix $\mu(\cdot)\in C^{\alpha}([0,T];\mathcal{P}_{1})$, $\alpha\in(0,1)$, and we consider the problem of choosing $(s(\cdot),v(\cdot))\in\mathcal{K}$ as to maximise $J(t_{0},x_{0},h_{0};v(\cdot),s(\cdot),\mu(\cdot))$ subject to (2) and (3), that is, to
@@ 3.1.2 Properties of the value function
$\begin{cases}-\partial_{t}V+\rho V=H_{1}(t,x,h,\mu(t),D_{h}V)+\frac{1}{2}h^{2}\chi^{2}D^{2}_{hh}V+\frac{1}{2}\epsilon^{2}D^{2}_{xx}V+H_{0}(D_{x}V)&\mbox{ in}(0,T)\times\mathbb{R}\times\mathbb{R}_{++}\\ V(T,x,h)=0&\mbox{ in }\mathbb{R}\times\mathbb{R}_{+}.\end{cases}$ (10)
@@
    $V(t,h,0)=0$ for every $(t,x)\in[0,T]\times\mathbb{R}$, and $V$ is strictly positive on $[0,T]\times\mathbb{R}\times\mathbb{R}_{++}$.
@@
    If, in addition, $\mu(\cdot)$ is uniformly Hölder continuous on [0,])[0,])[ 0 , ] ), then $V\in C^{1,2}((0,T)\times\mathbb{R}\times\mathbb{R}_{++})$.
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\right]\mathrm{d}t\\ \leq\frac{2^{\eta}}{1-\sigma}L_{f}^{\eta}e^{\frac{\eta}{2}C_{2,2}(T-t_{0})}(T-t_{0})h_{1}^{\eta}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right).$ (19)
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,2}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\mathrm{d}t\right]\\ \leq\frac{1}{1-\sigma}\overline{A}^{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}(T-t_{0})}(T-t_{0})\left(\left|h_{1}-h_{2}\right|^{2}+32TC_{1,2}L_{f}^{2}L_{\eta}^{2}h_{2}^{2}e^{C_{2}(T-t_{0})}\left|x_{1}-x_{2}\right|^{2}\right)^{\frac{\eta}{2}}.$ (20)
@@
    $V(t_{1},x_{0},h_{0})-V(t_{2},x_{\varepsilon}(t_{2}),h_{\varepsilon}(t_{2}))$ $\leq\mathbb{E}\left[\int_{t_{1}}^{t_{2}}e^{-\rho t}U_{\sigma}(x_{\varepsilon}(t),h_{\varepsilon}(t),v_{\varepsilon}(t),s_{\varepsilon}(t),\mu(t))\,\mathrm{d}t\right]+\varepsilon,$
@@
    $\cdot\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{2}\right]^{\frac{\gamma(1-\sigma)}{2}}+L_{A}^{1-\sigma}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{2}\right]^{\frac{1-\sigma}{2}}\right)$
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
    $\cdot\left(\mathbb{E}\left[\left|\tilde{h}-h_{0}\right|^{2}\right]+64TC_{1,2}L_{f}^{2}L_{\eta}^{2}e^{\frac{1}{2}C_{2,4}T}h_{0}^{2}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{4}\right]^{\frac{1}{2}}\right)^{\frac{\eta}{2}}$
@@
    $\cdot\left(\overline{A}^{1-\sigma}\left(\sqrt{8}\sqrt{K^{2}T+\epsilon^{2}}\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}(t_{2}-t_{1})^{\frac{\gamma(1-\sigma)}{2}}+\left(\sqrt{2}\sqrt{K^{2}T+\epsilon^{2}}L_{A}\right)^{1-\sigma}(t_{2}-t_{1})^{\frac{1-\sigma}{2}}\right)$
    $+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
    $\cdot\left(12C_{2,2}e^{C_{2,2}T}h_{0}^{2}(t_{2}-t_{1})+64TC_{1,2}L_{f}^{2}L_{\eta}^{2}e^{\frac{1}{2}C_{2,4}T}h_{0}^{2}\sqrt{8}\sqrt{T^{2}K^{4}+\epsilon^{4}}(t_{2}-t_{1})\right)^{\frac{\eta}{2}}.$
@@
    $V(t_{2},x_{0},h_{0})-V(t_{1},x_{0},h_{0})\leq\delta+\varepsilon+\widetilde{C}_{1}(t_{2}-t_{1})+\widetilde{C}_{2}(t_{2}-t_{1})^{\frac{\gamma(1-\sigma)}{2}}+\widetilde{C}_{3}(t_{2}-t_{1})^{\frac{1-\sigma}{2}}+\widetilde{C}_{4}(t_{2}-t_{1})^{\frac{\eta}{2}};$
@@
    Let $t_{0}\in\mathbb{R}_{++},x_{0}\in\mathbb{R},h_{0}\in\mathbb{R}_{++}$ and take first $\varepsilon>0$ such that $h_{0}-\varepsilon\in\mathbb{R}_{++}$. Define

### v3

@@ Abstract
We study a family of mean field games arising in modeling the behavior of strategic economic agents which move across space maximizing their utility from consumption and have the possibility to accumulate resources for production (such as human capital). The resulting mean field game PDE system is not covered in the actual literature on the topic as it displays weaker assumptions on the regularity of the data (in particular global Lipschitz continuity and boundedness of the objective are lost), state constraints, and a non-standard interaction term. We obtain a first result on the existence of solution of the mean field game PDE system.
@@
Daria Ghilli has been supported by the INdAM-GNAMPA project “Modelli Matematici per i Processi Decisionali riguardanti la Transizione Energetica” CUP $E53C22001930001$ and by the INdAM-GNAMPA project ”Modelli MFGs in Economia per lo studio della dinamica del capitale umano con spillovers spaziali” CUP $E55F22000270001.$ Giovanni Zanco has been supported by the F-NF grant “MFGPRE: Mean-field games: regularity, path-dependency, economics”.
@@ 1 Introduction
The departure point of the present paper is the problem of understanding the time-space evolution of an economic system, which is a crucial topic for economists and policy makers. The inclusion of the space dimension in economic analysis has regained relevance in the recent years. The emergence of a new economic geography is indeed one of the major events in the economic literature of the last decade (see e.g. [17], [18], [14]). Many papers have studied this problem under the assumption that decision about the system are taken by a unique agent (the so-called ”social planner”, which can be seen as the government of a nation): in this case the resulting mathematical problem is the optimal control of a Partial Differential Equation (PDE from now on), possibly stochastic (see e.g. [4, 3, 6, 15]).
@@
For a large population of agents and from the methodological viewpoint, the collective behavior of such a type of agents is effectively modeled using a Mean Field Game (MFG hereafter) framework. This approach heuristically emerges as the number of agents approaches infinity, condensing the complex multi-agent interactions into a tractable system of two coupled partial differential equations, a Hamilton-Jacobi-Bellman (HJB) equation, which captures the optimal control strategy of a representative agent and a Fokker-Planck (FP) equation, which describes the evolving spatial and human capital distribution of the entire population. The theory of MFGs has been intensively used in the last 20 years, starting from the seminal papers of Lasry-Lions [20] and Huang-Caines-Malhamé [16], and significantly allows to shed some light on the outcome of such complex dynamics. Moreover, the MFG methodology is motivated by two principal advantages. First, solutions to the MFG system are known to approximate the Nash equilibria of the discrete, finite-agent model with high accuracy for large N. Second, it offers a crucial computational benefit. Directly solving the N-coupled optimization problems of the discrete model quickly becomes intractable, even for a modest number of agents (e.g., $N=10$). In contrast, the MFG formulation reduces the problem to a system of just the HJB and FP equations, making numerical simulation and analysis far more feasible. This theory has been successfully employed to study the behavior of some economic models where many small homogeneous agents interact, see e.g. [1]. However, no paper tried to consider the case of the time-space evolution of economic variables, maybe because it displays many technical difficulties to which we return below in this introduction. We mention here the paper [12] where a kind of mean field approach is used, but without assuming that the small agents are forward-looking.
Hence, the purpose of this paper is to provide a first mathematical and theoretical analysis of a family of MFGs arising in modeling how the strategic interactions of small agents can shape the time-space evolution of economic variables. Further research will be subsequently devoted to the economic implications of such mathematical results, and numerical analysis of the MFG under consideration in the present paper will be carried out. We mention the paper [11] where such numerical analysis has already been used in order to study the evolution of a spatial epidemiological model where the rate of infection depends on the distribution of the agents.
@@
As a starting point, we assume that the small agents are homogeneous, move across space maximizing their intertemporal utility from consumption minus a disutility from displacement, and have the possibility to accumulate resources for production, for example, human capital. Hence their state variables, at any time $t\geq 0$, are the position $x(t)$ and the human capital $h(t)$. The control variables are the velocity $v(t)$ with which they move in the space, and the allocation of human capital production $s(t)$. We observe that the spatial and temporal progression of human capital is a subject of growing significance in economics. We refer e.g. to [22],one of the most important references for modeling human capital as an engine of growth, to [5] establishing an existing literature on human capital in an epidemic setting and to [2] for empirical evidence. A key feature of the model we propose is the presence of a spatial interaction terms influencing both the dynamics of the human capital and the utility of the agents. In economics such terms are called spatial spillovers, meaning that an individual’s human capital is dynamically influenced by that of their neighbors.
@@
The resulting MFG, even in the simplest case, turns out to be very difficult and quite far from what is covered in the actual literature on the topic (see e.g. the books [9, 7]). The main reasons are the presence of the above mentioned non standard interaction terms, the weaker regularity assumptions on the data-in particular Lipschitz continuity and boundedness of the objective are lost, and, finally, a state constraint structure since human capital needs to remain positive. In particular, the interaction term-which we will denote by $F$- has a non trivial mathematical structure, that is, is not globally Lipschitz with respect to the distribution. This feature (see Lemma 3.1 3.) is stricly connected to the fact that F models the spatial spillovers effect, and in particular that aggregation of agents increases the human capital $h$. Moreover, the positivity state constraints on the human capital brings the need of treating carefully the HJB and the FP equation proving suitable estimates near the border where $h=0$.
We study such MFG using the PDE approach, i.e. we consider the associated PDE system of Hamilton-Jacobi-Bellman (HJB) and Fokker-Planck (FP) equations. An intensive work has been carried out on such equations, here we mention only the ones which are more related to our setting. In particular some work on Fokker-Planck equations for uncontrolled interacting systems with spatial structure and non-standard interaction term has been carried out in [13] and [24]. Our main result is Theorem 4.4 on the existence of solutions of such PDE system. This is a departure point to understand the equilibria of the economic system, which will be the object of future research. In order to achieve this result we first study the HJB equation (see Proposition 3.3). Concerning the FP equation we first analyse the properties of the drift of such equation and in Lemma 3.6 we prove that such drift is locally Lispchitz continuous for a suitable set of distributions. These results are based on careful estimates of the underlying stochastic process modelling human capital (see e.g. point 3. in Proposition 3.3). After we turn to the FP equation and in order to establish well-posedness of such equation we analyse first the associated McKean-Vlasov equation. Existence of a strong solution of such equation, together with her pathwise uniqueness, are proved in Proposition 3.7. The proof is based on the estimate of the drift of the FP equation mentioned before, on a careful study of the properties of the underlying stochastic processes and on a fixed point argument. Once this result is carried out, we are ready to study the MFG in 4.4 through again a fixed point argument. The proof is based on the results on the HJB and FP already explained. We underline that our main existence result is local in time: in Theorem 4.4 we require the time horizon $T$ to belong to the admissible interval (see (41)). This restriction was already fundamental to ensure well-posedness of the FP equation via the associated McKean-Vlasov dynamics, see Proposition 3.7. This small-time condition comes from the a priori estimates in the FP and McKean-Vlasov analysis, which rely on Gronwall-type arguments and involve constants that grow exponentially in $T$. Hence we need to choose $T$ sufficiently small in order to close the fixed-point/invariance estimates.
@@ 2.1 The economic model
The state equations of each agent on time intervals $t_{0},T]$ are as follows. The dynamics for the position is given by
@@
[…]
@@
From the definition of F is straightworward to notice that F models the spatial interactions and the aggregation effect which increases the human capital of one individual. In other word,s the level of human capital of an individual is influenced bt the level of human capital of neighbourhoods. The” level” of influence is described mathematically by the functions $\eta_{i}$. As an example, we could consider $\eta_{i}$ being suitable regularizations of the indicator function of two balls, whose radius model the maximal distances of the heighbourhoods who influence the human capital of the agent considered.
@@
Individuals control their position in space to maximise their gain from such spillover, considering at the same time that moving towards areas with high human capital is a costly action. Assuming that no individual has an overwhelming influence on the systemwith respect to others, optimisation of each individual’s position gives rise to a symmetric game among economic agents (the individuals).
@@
$J(t_{0},x_{0},h_{0};v(\cdot),s(\cdot))\\ :=\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}(u_{\sigma}\left([(1-s(t))f(h(t))]^{1-\gamma}F(x(t),\mu(t))^{\gamma}A(x(t))\right)-a(v(t)))dt\right]$ (4)
over all admissible control process $(v,s)\in\mathcal{K}$. Here $\rho>0$ is the intertemporal discount factor while $a\colon K\rightarrow\mathbb{R}$ is a given strictly convex cost function satisfying $a(0)=0$. Moreover, A represents in economics the production function or amenity of local services.
We make the following further assumptions:
@@ 2.2 The Mean Field Game system
On $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ we consider the forward-backward system of PDEs
[…]
whose solution is a couple $(V,\mu)$, where $V$ is a real-valued function on $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ and $\mu$ is a function on $[0,T]$ taking values in $\mathcal{P}_{2}$ (a precise notion of solution will be given in Definition 2.3 below).
@@
    $V$ is continuous on $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$;
@@
    $D_{x}V$ and $D_{h}V$ are defined on $0,T]\times\mathbb{R}\times\mathbb{R}_{+}$;
@@
    $\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(t,x,h)$ $\mu(t;\mathrm{d}x,\mathrm{d}h)-\int_{\mathbb{R}\times\mathbb{R}_{+}}\phi(0,x,h)\mu_{0}(\mathrm{d}x,\mathrm{d}h)$ (7)
@@
[…]
@@
In particular we will often deal with maps $0,T]\ni t\mapsto\mu(t)$; in this case we write
@@ 3.1 The Hamilton-Jacobi-Bellman equation
To study the HJB equation we fix a family of measures $\mu(\cdot)\in C^{\alpha}([0,T];\mathcal{P}_{1})$, $\alpha\in(0,1)$, and we consider the problem of choosing $(s(\cdot),v(\cdot))\in\mathcal{K}$ as to maximise $J(t_{0},x_{0},h_{0};v(\cdot),s(\cdot),\mu(\cdot))$ subject to (2) and (3), that is, to
@@
Remark 3.2.
For fixed $x$, the function $\mu\mapsto F(x,\mu)$ is in general only locally Lipschitz continuous, not globally. Indeed $F$ is differentiable in the sense of Lions (at least for $y\neq x$) and its gradient has first component of the form $\partial_{\mu}F(x,\mu)(y,k)=C_{1}(x-y)k-C_{2}(x-y)\int C_{3}(x-y)k\mu(\mathrm{d}y,\mathrm{d}k)$ with $C_{1},C_{2},C_{3}$ bounded functions; the quantity $\int\partial_{\mu}F(x,\mu)(y,k)\mu(\mathrm{d}y,\mathrm{d}k)$ is unbounded over $\mathcal{P}_{2}$, whereas it should be bounded if $F$ were globally Lipschitz (see e.g. [9, Vol. I, Chapter 5]).
@@ 3.1.2 Properties of the value function
[…]
@@
    $V(t,h,0)=0$ for every $(t,x)\in[0,T]\times\mathbb{R}$, and $V$ is strictly positive on $0,T]\times\mathbb{R}\times\mathbb{R}_{++}$.
@@
    If, in addition, $\mu(\cdot)$ is uniformly Hölder continuous on $[0,T)$, then $V\in C^{1,2}((0,T)\times\mathbb{R}\times\mathbb{R}_{++})$.
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,2}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\right]\mathrm{d}t\\ \leq\frac{2^{\eta}}{1-\sigma}L_{f}^{\eta}e^{\frac{\eta}{2}C_{2,2}(T-t_{0})}(T-t_{0})h_{1}^{\eta}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\\ \cdot\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\left|x_{1}-x_{2}\right|^{\gamma(1-\sigma)}+L_{A}^{1-\sigma}\left|x_{1}-x_{2}\right|^{1-\sigma}\right).$ (19)
@@
    $\mathbb{E}\left[\int_{t_{0}}^{T}e^{-\rho t}\left|U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,1}(t),s_{\varepsilon}(t),\mu(t)\right)-U_{\sigma}\left(x_{\varepsilon,1}(t),h_{\varepsilon,2}(t),s_{\varepsilon}(t),\mu(t)\right)\right|\mathrm{d}t\right]\\ \leq\frac{1}{1-\sigma}\overline{A}^{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu(\cdot))\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}(T-t_{0})}(T-t_{0})\\ \cdot\left(\left|h_{1}-h_{2}\right|^{2}+32TC_{1,2}L_{f}^{2}L_{\eta}^{2}h_{2}^{2}e^{C_{2}(T-t_{0})}\left|x_{1}-x_{2}\right|^{2}\right)^{\frac{\eta}{2}}.$ (20)
@@
    $V(t_{1},x_{0},h_{0})-V(t_{2},x_{\varepsilon}(t_{2}),$ $h_{\varepsilon}(t_{2}))$
    $\leq\mathbb{E}\left[\int_{t_{1}}^{t_{2}}e^{-\rho t}U_{\sigma}(x_{\varepsilon}(t),h_{\varepsilon}(t),v_{\varepsilon}(t),s_{\varepsilon}(t),\mu(t))\,\mathrm{d}t\right]+\varepsilon,$
@@
    $\phantom{aa}\cdot\left(\overline{A}^{1-\sigma}\left(2\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{2}\right]^{\frac{\gamma(1-\sigma)}{2}}+L_{A}^{1-\sigma}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{2}\right]^{\frac{1-\sigma}{2}}\right)$
    $\phantom{aa}+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
    $\phantom{aa}\cdot\left(\mathbb{E}\left[\left|\tilde{h}-h_{0}\right|^{2}\right]+64TC_{1,2}L_{f}^{2}L_{\eta}^{2}e^{\frac{1}{2}C_{2,4}T}h_{0}^{2}\mathbb{E}\left[\left|\tilde{x}-x_{0}\right|^{4}\right]^{\frac{1}{2}}\right)^{\frac{\eta}{2}}$
@@
    $\phantom{aa}\cdot\left(\overline{A}^{1-\sigma}\left(\sqrt{8}\sqrt{K^{2}T+\epsilon^{2}}\frac{L_{\eta}}{\theta}\right)^{\gamma(1-\sigma)}(t_{2}-t_{1})^{\frac{\gamma(1-\sigma)}{2}}+\left(\sqrt{2}\sqrt{K^{2}T+\epsilon^{2}}L_{A}\right)^{1-\sigma}(t_{2}-t_{1})^{\frac{1-\sigma}{2}}\right)$
    $\phantom{aa}+\frac{\overline{A}^{1-\sigma}}{1-\sigma}\left(\frac{\Theta}{\theta}\overline{M}(\mu)\right)^{\gamma(1-\sigma)}\left(2L_{f}\right)^{\eta}e^{\eta C_{2,2}T}T$
    $\phantom{aaa}\cdot\left(12C_{2,2}e^{C_{2,2}T}h_{0}^{2}(t_{2}-t_{1})+64TC_{1,2}L_{f}^{2}L_{\eta}^{2}e^{\frac{1}{2}C_{2,4}T}h_{0}^{2}\sqrt{8}\sqrt{T^{2}K^{4}+\epsilon^{4}}(t_{2}-t_{1})\right)^{\frac{\eta}{2}}.$
@@
    $V(t_{2},x_{0},h_{0})-V(t_{1},x_{0},h_{0})\\ \leq\delta+\varepsilon+\widetilde{C}_{1}(t_{2}-t_{1})+\widetilde{C}_{2}(t_{2}-t_{1})^{\frac{\gamma(1-\sigma)}{2}}+\widetilde{C}_{3}(t_{2}-t_{1})^{\frac{1-\sigma}{2}}+\widetilde{C}_{4}(t_{2}-t_{1})^{\frac{\eta}{2}};$
@@
    Let $t_{0}\in\mathbb{R}_{++},x_{0}\in\mathbb{R},h_{0}\in\mathbb{R}_{++}$ and take first $\varepsilon>0$ such that $h_{0}-\varepsilon\in\mathbb{R}_{++}$ and $(t_{0}-\varepsilon,t_{0}+\varepsilon)\in[0,T)$. Define